Mathematics 241 – Multivariable Calculus Quiz 6 – November 2, 2007

Your Name:

A) (5) Find the curl of the vector field F = (cos(x)e^{y}, sin(y)e^{z}, cos(z)e^{x}).

B) (5) Find second-degree Taylor polynomial of f(x, y) = xe^{−y^{2}} at (x, y) = (0, 1)